OpenStudy (anonymous):

(s-3)/(s^2-6s+25) inverse laplace transform

OpenStudy (anonymous):

hint : the denom can be re-written as (s^2-6s+25) = (s^2-6s+9+16)=(s^2-6s+9)+ 16 =(s-3)^2+4^2

OpenStudy (anonymous):

AH NICE thankyou.... got it. ihave 6 more question help me please thanks..

OpenStudy (anonymous):

(e^3s)/(s+3) inverse laplace